OpenStudy (anonymous):

X - 7 = -3 Solve for the variable. Write the answer only, do not type in X= form

OpenStudy (anonymous):

@theredhead1617 substitute the -7 Can you do that?

OpenStudy (anonymous):

Subsitute it with what?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

just like the other one

OpenStudy (anonymous):

So the answer is 4?

OpenStudy (anonymous):

no substitution needed @Trexy X -7 = -3 X = -3 + 7 X = 4

OpenStudy (anonymous):

LOL @miteshchvm sorry i keep saying substitute.. I've been doin math homework too, @theredhead1617 Its not Substitute SORRY its my fault  just change the sign and put it to the other side. xD